Please help me answer this problem.I really dont know
the formula.
Will had an average of 74 on his first four math test.after taking the next test,his average dropped to 73.find his most recent test grade

Answer by rothauserc(4718)

need to use the formula for average
average = sum of test scores / number of test scores
73 = (4 *74 + recent test score) / 5
365 = 296 + recent test score
the recent test score = 365 - 296 = 69